OpenStudy (anonymous):

Find the y intercept of the parabola defined by y=2x^2+x-1

OpenStudy (amistre64):

y intercept means "x=0" what value do you get when x=0?

OpenStudy (anonymous):

hang on

OpenStudy (amistre64):

think about it like this; everthing with an x in it goes away :)

OpenStudy (anonymous):

....so just -1? idk..

OpenStudy (amistre64):

i believe so; lets compare: y=2x^2+x -1 y = 0 + 0 -1 y = -1 yep

OpenStudy (anonymous):

but 0,-1? I have to have an ordered pair.

OpenStudy (amistre64):

when (x=0,y=-1) ( 0, -1); or just (0,-1)